Express
(2,π/6 in rectangular coordinates

Express r=1+2cos∅ from polar to cartesian form
1
Expert's answer
2020-08-27T16:00:04-0400

I)Express "(2,\frac{\pi}{6})" in rectangular coordinates:

"x=rcos\theta,"

"y=rsin\theta;"

"r=2,\space \theta=\frac{\pi}{6}"

"x=2*cos(\frac{\pi}{6}),"

"y=2*sin(\frac{\pi}{6})."

"x=2*\frac{\sqrt{3}}{2}=\sqrt{3},"

"y=2*\frac{1}{2}=1."

Answer:"x=\sqrt{3},\space y=1."

II)Express "r=1+2cos\theta" from polar to cartesian form.

"x=rcos\theta,"

"y=sin\theta."

"r=\sqrt{x^2+y^2}"

"r=1+2cos\theta"

"2cos\theta=r-1"

"2(\frac{x}{r})=r-1"

"2x=r^2-r"

"2x=(x^2+y^2)-\sqrt{x^2+y^2}"

Answer:

"x^2+y^2=2x+\sqrt{x^2+y^2}."